# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v5

@@ 3.2 Privacy blanket framework
The decomposition of the standard clone is constructed by treating $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ as reference points and projecting each $\mathcal{R}(x_{i})$ onto the basis defined by these two distributions [FMT22]. In contrast, the decomposition provided by the privacy blanket framework takes a different approach: it first identifies the “common part” shared across all $\mathcal{R}(x)$ distributions [BBGN19]:
@@
Here, $\omega$ and $\gamma$ are referred to as the privacy blanket distribution and the total variation similarity of the local randomizer $\mathcal{R}$ [BBGN19]. Each $\mathcal{R}(x)$ can then be decomposed as:
@@
The original proof is formulated using the terminology of the “View” of the server. We restate the privacy blanket technique using the following notation: probability distributions $P_{0}^{\mathsf{B}}$ and $P_{1}^{\mathsf{B}}$ (with “$\mathsf{B}$” denoting “Blanket”), along with a post-processing function $f^{\mathsf{B}}$.
Definition 6 (Privacy Blanket Reduction Pair [BBGN19] (Restated)).
Let $\boldsymbol{x}_{-1}=(x_{2},x_{3},\dots,x_{n})$ with the inputs from the last $n-1$ users, $\boldsymbol{y}^{a}=(y_{1}^{a},y_{2},\dots,y_{n})$ where $y_{i}\sim R(x_{i})$ is the output of the i-th user, $a$ indicates that the input of the first user is $x_{1}^{a}$, $a\in\{0,1\}$. Let $\boldsymbol{b}=(b_{2},b_{3},\dots,b_{n})$ be binary values indicating which users sample from the privacy blanket distribution. A multiset $Y_{\boldsymbol{b}}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|b_{i}=1\})$.
Observe that the distribution of $Y_{\boldsymbol{b}}^{a}$ depends only on $|\boldsymbol{b}|$ rather than $\boldsymbol{b}$, where $|\boldsymbol{b}|$ represents the number of 1 in $\boldsymbol{b}$. We can rewrite it as $Y_{|\boldsymbol{b}|}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|y_{i}\sim\omega,i=1,2,\dots,|\boldsymbol{b}|\})$. Then $P_{0}^{\mathsf{B}}$ and $P_{1}^{\mathsf{B}}$ are defined below:
@@
Theorem 2 (Privacy Blanket Reduction [BBGN19] (Restated)).
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}}\big{)}.$
@@
$x_{2},\dots,x_{n}$; $|\boldsymbol{b}|\in\{0,1,\dots,n-1\},Y^{a}_{|\boldsymbol{b}|}$
@@
     Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
     $s\leftarrow_{\text{LO}(x_{j})}\mathbb{Y}$triangleright Sample from $\text{LO}(x_{j})$
@@ 3.3 Vision and failure of stronger clone
The stronger clone is expected to improve the probability of producing a “clone” from $\frac{1}{2e^{\varepsilon_{0}}}$ to $\frac{1}{e^{\varepsilon_{0}}+1}$. For $\varepsilon_{0}>1$, this results in approximately a factor of 2 improvement in the expected number of “clones” [FMT23a]. This improvement is anticipated to be achieved through a more refined analysis that, instead of cloning the entire output distributions on differing elements, clones only the portions of those distributions where they actually differ.
Specifically, it leverages a lemma from [YB17] to establish the existence of the following decomposition:
Theorem 3 (Corollary 3.4 in [FMT23a]).
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\parallel P_{1}(\varepsilon_{0},p)\big{)},$
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1}),P_{1}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1})\big{)}.$ (1)
Lemma 2 (Lemma 5.1. in [FMT23a]).
@@
$D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p)\parallel P_{1}(\varepsilon_{0},p)\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p^{\prime})\parallel P_{1}(\varepsilon_{0},p^{\prime})\big{)}$
Unfortunately, they encountered difficulty in constructing a post-processing function $f$ for this construction of $P_{0}$ and $P_{1}$. While they provided a function in the original paper, it was proven to be incorrect in the corrected revision [FMT23b]. The issue arises from the fact that the “leftover” distribution of $x_{1}$ (i.e., $\mathcal{Q}_{1}$) is mixed with the “leftover” distribution of $x_{i}$ (i.e., $\mathcal{Q}_{i}$) in the above construction. In this case, the function $f$ does not know which distribution to sample from. This technical problem is fundamental and remains unsolved.
Although the corrected version was published on arXiv in October 2023, this error has been propagated in subsequent works [WPL+24, CCG24, YYF+24, WZL+25, KHH23]. For instance, the variation-ratio framework made significant efforts to design an algorithm to find the parameter $p$ for various specific local randomizers in the decomposition of Theorem 3. However, their work relies on the incorrect post-processing function $f$ presented in [FMT23a], which renders their results invalid.
Due to this fundamental difficulty, it is required that $x_{1}$ has no “leftover” distribution. In other words, each component of $\mathcal{R}(x_{1})$ must be distinguishable from the “leftover” distribution of $x_{i}$ for $i\geq 2$. In the above example, this necessitates a four-point-based construction for $P_{0}(\varepsilon_{0},p,q)$ and $P_{1}(\varepsilon_{0},p,q)$ [FMT23b]:
@@
This new decomposition proposed in the corrected version introduces additional challenges. First, for specific randomizers, computing a tight value of $q$ is nontrivial. Second, the monotonicity of $D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p,q)\parallel P_{1}(\varepsilon_{0},p,q)\big{)}$ with respect to $p$ is not known. Consequently, we are unable to derive the desired conclusion—namely, a general upper bound applicable to any $\varepsilon_{0}$-DP local randomizer, as stated in (1). For the same reason, it remains unclear whether this decomposition necessarily yields tighter bounds than the standard clone decomposition. Moreover, the new bound $D_{e^{\varepsilon}}\big{(}P_{0}(\varepsilon_{0},p,q)\parallel P_{1}(\varepsilon_{0},p,q)\big{)}$ for a specific local randomizer lacks an efficient algorithm to compute.
@@ 4 General Clone and the Optimal Bounds
Figure 1: Hierarchy among decomposition-based methods
@@ 4.1 Definition of general clone
The reduction pair $(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$ (with “$\mathsf{GC}$” denoting “General Clone”) is defined as the distributions of the histograms over $\{1,\dots,k+1\}$ generated by sampling:
@@
$D_{e^{\varepsilon}}\big{(}\mathcal{M}_{S}(X^{0})\parallel\mathcal{M}_{S}(X^{1})\big{)}\leq D_{e^{\varepsilon}}\big{(}P_{0}^{\mathsf{GC}}\parallel P_{1}^{\mathsf{GC}}\big{)}.$
@@
The corresponding post-processing function $f^{\mathsf{GC}}$ is shown in Algorithm 3. It operates by iterating over the input. When encountering indices $1,2,\dots,k$, it samples once from each corresponding distribution $Q_{1}^{1},Q_{1}^{2},\dots,Q_{1}^{k}$. When encountering index $k+1$, it uniformly selects an element $j\in\{2,3,\dots,n\}$ that has not yet been selected, and samples once from $Q_{j}$, which is the left-over distribution corresponding to $x_{j}$. Finally, it returns all the sampled values as output.
We now prove that $f^{\mathsf{GC}}(P_{0}^{\mathsf{GC}})=\mathcal{M}_{S}(X^{0})$; the proof for $f^{\mathsf{GC}}(P_{1}^{\mathsf{GC}})=\mathcal{M}_{S}(X^{1})$ is analogous.
Recall that $P_{0}^{\mathsf{GC}}$ is a histogram over a mixture of one sample from $A_{0}$ and $n-1$ samples from $A_{2}$, while $\mathcal{M}_{S}(X^{0})$ is the shuffled mixture of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{i})$ for $i=2,3,\dots,n$. A well-known property is that producing a histogram over outputs is equivalent to applying a random shuffling to the output multiset—each can be generated from the other [CSWS21]. For clarity of exposition, we adopt the latter interpretation.
Let $\mathcal{P}_{n}$ denote the set of all permutations over $[n]$. In $P_{0}^{\mathsf{GC}}$, the sampling is followed by a uniformly random permutation $\sigma_{1}\leftarrow_{\$}\mathcal{P}_{n}$, such that the $\sigma_{1}(1)$-th element $z_{\sigma_{1}(1)}$ is drawn from $A_{0}$. The remaining positions $\sigma_{1}(i)$, for $i=2,3,\dots,n$, are drawn from $A_{2}$. It can be justified that $f^{\mathsf{GC}}(z_{\sigma_{1}(1)})\sim\mathcal{R}(x_{1}^{0})$:
$\Pr[f^{\mathsf{GC}}(z_{\sigma_{1}(1)})=y]=\Pr[f^{\mathsf{GC}}(A_{0})=y]$ $=\sum_{m=1}^{k}\Pr[A_{0}=m]\cdot\Pr[f^{\mathsf{GC}}(m)=y]$
$=\sum_{j=1}^{k}a_{j}Q_{1}^{j}(y)=\mathcal{R}(x_{1}^{0})(y).$
An important observation is that the support of $A_{0}$ does not include the symbol $k+1$. As a result, the on-the-fly processing of $k+1$ in $f^{\mathsf{GC}}$ is equivalent to introducing a new permutation $\sigma_{2}\leftarrow_{\$}\{\sigma\in\mathcal{P}_{n}\mid\sigma(\sigma_{1}(1))=\sigma_{1}(1)\}$. Under this view, if $z_{i}=k+1$, then sampling from $Q_{\sigma_{2}(\sigma_{1}(i))}$ is equivalent to the original procedure.
We can then show that the output of $f^{\mathsf{GC}}$ at round $i$ (denoted by $f_{i}^{\mathsf{GC}}(z_{i})$ ) satisfies $f_{i}^{\mathsf{GC}}(z_{i})\sim\mathcal{R}(x_{\sigma_{2}(\sigma_{1}(i))})$, for all $i\neq\sigma_{1}(1)$:
$\Pr[f^{\mathsf{GC}}_{i}(z_{i})=y]$ $=\Pr[f^{\mathsf{GC}}_{i}(A_{2})=y]$
$=\sum_{m=1}^{k}\Pr[A_{2}=m]\cdot\Pr[f^{\mathsf{GC}}_{i}(m)=y]+\Pr[A_{2}=k+1]\cdot\Pr[f_{i}^{GC}(k+1)=y]$
$=\sum_{m=1}^{k}c_{j}Q_{1}^{m}(y)+\beta Q_{\sigma_{2}(\sigma_{1}(i))}(y)$
$=\mathcal{R}(x_{\sigma_{2}(\sigma_{1}(i))}).$
It can be shown that the composition $\sigma_{2}\circ\sigma_{1}$ is also uniformly distributed over $\mathcal{P}_{n}$. In summary, the output of $f^{\mathsf{GC}}(P_{0}^{\mathsf{GC}})$ is a random shuffling over the multiset consisting of $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{i})$ for $i=2,3,\dots,n$, which completes the proof. ∎
@@
$x_{1}^{0},x_{1}^{1},x_{2},\dots,x_{n}$; $z\in\{1,2,\dots,k+1\}^{n}$
$J\leftarrow\emptyset$
for $i=1,\dots,n$ do
     if $z_{i}=k+1$ then
         Let $j$ be a randomly and uniformly chosen element of $[2:n]\setminus J$
         $J\leftarrow J\cup\{j\}$
@@
     Sample $y_{i}$ from
      $\begin{cases}Q_{1}^{z_{i}}&\text{if }z_{i}\in\{1,2,\dots,k\};\\ Q_{j}&\text{if }z_{i}=k+1.\end{cases}$triangleright Sample from the left-over distribution of $x_{j}$ when $z_{i}=k+1$

### v6

@@ 3.2 Privacy blanket framework
The decomposition of the standard clone is constructed by treating $\mathcal{R}(x_{1}^{0})$ and $\mathcal{R}(x_{1}^{1})$ as reference points and projecting each $\mathcal{R}(x_{i})$ onto the basis defined by these two distributions [Feldman2021]. In contrast, the decomposition provided by the privacy blanket framework takes a different approach: it first identifies the “common part” shared across all $\mathcal{R}(x)$ distributions [Balle2019]:
@@
Here, $\omega$ and $\gamma$ are referred to as the privacy blanket distribution and the total variation similarity of the local randomizer $\mathcal{R}$ [Balle2019]. Each $\mathcal{R}(x)$ can then be decomposed as:
@@
The original proof is formulated using the terminology of the “View” of the server. We restate the privacy blanket technique using the following notation: probability distributions $P_{0}^{\mathsf{B}}$ and $P_{1}^{\mathsf{B}}$ (with $\mathsf{B}$” denoting “Blanket”), along with a post-processing function $f^{\mathsf{B}}$.
Definition 6 (Privacy Blanket Reduction Pair [Balle2019] (Restated)).
Let $\boldsymbol{x}_{-1}=(x_{2},x_{3},\dots,x_{n})$ with the inputs from the last n1n-1 users, $\boldsymbol{y}^{a}=(y_{1}^{a},y_{2},\dots,y_{n})$ where $y_{i}\sim R(x_{i})$ is the output of the i-th user, $a$ indicates that the input of the first user is $x_{1}^{a}$, $a\in\{0,1\}$. Let $\boldsymbol{b}=(b_{2},b_{3},\dots,b_{n})$ be binary values indicating which users sample from the privacy blanket distribution. A multiset $Y_{\boldsymbol{b}}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|b_{i}=1\})$.
Observe that the distribution of $Y_{\boldsymbol{b}}^{a}$ depends only on $|\boldsymbol{b}|$ rather than $\boldsymbol{b}$, where $|\boldsymbol{b}|$ represents the number of 1 in $\boldsymbol{b}$. We can rewrite it as $Y_{|\boldsymbol{b}|}^{a}=\mathcal{S}(\{y_{1}^{a}\}\cup\{y_{i}|y_{i}\sim\omega,i=1,2,\dots,|\boldsymbol{b}|\})$. Then $P_{0}^{B}$ and $P_{1}^{B}$ are defined below:
@@
Theorem 2 (Privacy Blanket Reduction [Balle2019] (Restated)).
@@
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq D_{e^{\varepsilon}}\big(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}}\big).$
@@
  Meta parameter: $x_{2},\dots,x_{n}$
$|\boldsymbol{b}|\in\{0,1,\dots,n-1\},Y^{a}_{|\boldsymbol{b}|}$
@@
  Let $j$ be a randomly and uniformly chosen element of $2:n]\setminus J$
  $s\leftarrow_{\text{LO}(x_{j})}\mathbb{Y}$ $\triangleright$ Sample from $\text{LO}(x_{j})$
@@ 3.3 Vision and failure of stronger clone
The stronger clone is expected to improve the probability of producing a “clone” from $\frac{1}{2e^{\varepsilon_{0}}}$ to $\frac{1}{e^{\varepsilon_{0}}+1}$. For $\varepsilon_{0}>1$, this results in approximately a factor of 2 improvement in the expected number of “clones” [feldman2023soda]. This improvement is anticipated to be achieved through a more refined analysis that, instead of cloning the entire output distributions on differing elements, clones only the portions of those distributions where they actually differ.
Specifically, it leverages a lemma from [Ye] to establish the existence of the following decomposition:
Theorem 3 (Corollary 3.4 in [feldman2023soda]).
@@
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},p),P_{1}(\varepsilon_{0},p)\big),$
@@
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1}),P_{1}(\varepsilon_{0},\frac{1}{e^{\varepsilon_{0}}+1})\big).$ (1)
Lemma 2 (Lemma 5.1. in [feldman2023soda]).
@@
$D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},p)\,\|\,P_{1}(\varepsilon_{0},p)\big)\leq D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},p^{\prime})\,\|\,P_{1}(\varepsilon_{0},p^{\prime})\big)$
Unfortunately, they encountered difficulty in constructing a post-processing function $f$ for this construction of $P_{0}$ and $P_{1}$. While they provided a function in the original paper, it was proven to be incorrect in the corrected revision [feldman2023arxiv]. The issue arises from the fact that the “leftover” distribution of $x_{1}$ (i.e., $\mathcal{Q}_{1}$) is mixed with the “leftover” distribution of $x_{i}$ (i.e., $\mathcal{Q}_{i}$) in the above construction. In this case, the function $f$ does not know which distribution to sample from. This technical problem is fundamental and remains unsolved.
Although the corrected version was published on arXiv in October 2023, this error has been propagated in subsequent works [tv_shuffle, ChenCG24, wang1, wang2, koskela2023numerical]. For instance, the variation-ratio framework made significant efforts to design an algorithm to find the parameter $p$ for various specific local randomizers in the decomposition of Theorem 3. However, their work relies on the incorrect post-processing function $f$ presented in [feldman2023soda], which renders their results invalid.
Due to this fundamental difficulty, it is required that $x_{1}$ has no “leftover” distribution. In other words, each component of $\mathcal{R}(x_{1})$ must be distinguishable from the “leftover” distribution of $x_{i}$ for $i\geq 2$. In the above example, this necessitates a four-point-based construction for $P_{0}(\varepsilon_{0},p,q)$ and $P_{1}(\varepsilon_{0},p,q)$ [feldman2023arxiv]:
@@
This new decomposition proposed in the corrected version introduces additional challenges. First, for specific randomizers, computing a tight value of $q$ is nontrivial. Second, the monotonicity of $D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},p,q),P_{1}(\varepsilon_{0},p,q)\big)$ with respect to $p$ is not known. Consequently, we are unable to derive the desired conclusion—namely, a general upper bound applicable to any $\varepsilon_{0}$-DP local randomizer, as stated in (1). For the same reason, it remains unclear whether this decomposition necessarily yields tighter bounds than the standard clone decomposition. Moreover, the new bound $D_{e^{\varepsilon}}\big(P_{0}(\varepsilon_{0},p,q),P_{1}(\varepsilon_{0},p,q)\big)$ for a specific local randomizer lacks an efficient algorithm to compute.
@@ 4 General Clone and the Optimal Bounds
Figure 1: Hierarchy among decomposition-based methods. The Standard clone is horizontal (generic bound), while the Blanket curve varies with the randomizer (specific bound).
@@ 4.1 Definition of general clone
The reduction pair $(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}})$ (with $\mathsf{GC}$” denoting “General Clone”) is defined as the distributions of the histograms over $\{1,\dots,k+1\}$ generated by sampling:
@@
$D_{e^{\varepsilon}}\big(\mathcal{M}_{S}(X^{0}),\mathcal{M}_{S}(X^{1})\big)\leq D_{e^{\varepsilon}}\big(P_{0}^{\mathsf{GC}},P_{1}^{\mathsf{GC}}\big).$
@@
The corresponding post-processing function $f^{\mathsf{GC}}$ is given in Algorithm 3. Its behavior is as follows. For each index $j\in\{1,2,\dots,k\}$, it draws $n_{j}$ samples from the distribution $Q_{1}^{j}$. For index $k+1$, it uniformly samples an $n_{k+1}$-subset $I\subseteq\{2,3,\dots,n\}$ and, for each $i\in I$, draws a sample from the “left-over” distribution $Q_{i}$. The algorithm finally outputs the multiset of all sampled values.
[…]
Recall that $P_{0}^{\mathsf{GC}}$ is the histogram of one sample drawn from $A_{0}$ and n1n-1 independent samples drawn from $A_{2}$, whereas $\mathcal{M}_{S}(X^{0})$ is obtained by shuffling the outputs of the local randomizer applied to $x_{1}^{0},x_{2},\dots,x_{n}$. By the general clone decomposition, $\mathcal{M}_{S}(X^{0})$ admits the following equivalent generative description: user 1 draws $a_{1}\sim A_{0}$ and then $y_{1}\sim Q_{1}^{a_{1}}$; each other user $i\geq 2$ draws $a_{i}\sim A_{2}$ and then $y_{i}\sim Q_{1}^{a_{i}}$ if $a_{i}\in[k]$, or $y_{i}\sim Q_{i}$ otherwise. The shuffled output is $\{y_{i}\}_{i=1}^{n}$.
Let $T^{a}=\{a_{i}\}$, $T^{y}=\{y_{i}\}$, and $T^{a,y}=\{(a_{i},y_{i})\}$ denote the latent, output, and joint multisets, respectively. Let $H^{a}=(n_{1},n_{2},\dots,n_{k+1})$ be the histogram of $T^{a}$, where $n_{j}=|\{i:a_{i}=j\}|$. Then the marginals of $H^{a}$ and $T^{y}$ are exactly $P_{0}^{\mathsf{GC}}$ and $\mathcal{M}_{S}(X^{0})$.
Define the index set $I=\{i:a_{i}=k+1\}.$ Conditioned on $H^{a}$, symmetry implies that $I$ is uniformly distributed over all subsets of $\{2,3,\dots,n\}$ of size $n_{k+1}$. Furthermore,
$T^{a,y}\mid(H^{a},I)$ $=\cup_{j=1}^{k}\{(j,y_{t}):\ y_{t}\sim Q_{1}^{j},t=1,2,\dots,n_{j}\}$
$\quad\cup\{(k+1,y_{i}):i\in I,\ y_{i}\sim Q_{i}\}.$
Thus, $T^{y}\mid(H^{a},I)=\Big(\bigcup_{j=1}^{k}V_{j}\Big)\cup V_{k+1},$
where $V_{j}$ consists of $n_{j}$ i.i.d. samples from $Q_{1}^{j}$ for $1\leq j\leq k$, and $V_{k+1}=\{y_{i}:i\in I,\ y_{i}\sim Q_{i}\}.$ This conditional generative process matches exactly the sampling performed by $f^{\mathsf{GC}}$. Therefore, for every histogram $h$, $f^{\mathsf{GC}}(h)\stackrel{{\scriptstyle d}}{{=}}T^{y}\mid(H^{a}=h),$ and hence $f^{\mathsf{GC}}(P_{0}^{\mathsf{GC}})\stackrel{{\scriptstyle d}}{{=}}\mathcal{M}_{S}(X^{0}).$
@@
$(n_{1},\dots,n_{k},n_{k+1})\in\{0,1,\dots,n\}^{k+1}$
$O\leftarrow\emptyset$
for $j=1,\dots,k+1$ do
  if $j=k+1$ then
   Let $I$ be a randomly and uniformly chosen $n_{k+1}$-size subset of $\{2,3,\dots,n\}$
   for $i\in I$ do
     $O\cup\{s\mid s\sim Q_{i}\}$ $\triangleright$ Sample from the left-over distribution of $x_{i}$
   end for
  else
   for $loop=1,2,\dots,n_{j}$ do
     $O\cup\{s\mid s\sim Q_{1}^{j}\}$ $\triangleright$ Sample from $Q^{j}_{1}$
   end for